i do not understand
f (x)= x^22 + 6x +9
o, -3

I need to know the steps to get the result

Answer by ewatrrr(24785)


Hi,
f (x)= x^2 + 6x +9
factored
f (x)= (x+3)(x+3)
|setting f(x) = 0 to find the root(s) of the equation
0 = (x+3)(x+3) |note factors are indentical is this example
(x+3)= 0
x = -3
PT(-3,0) is the x-intercept and the vertex of the parabola as well in this ex